# A Case Study on Misinformation Detection

### v3

@@ 1 INTRODUCTION
Online services face a daunting task: There is an unceasing deluge of user-generated content, on the order of hundreds of thousands of posts per minute on popular social media platforms [275]. Some of that content is false, hateful, harassing, extremist, or otherwise problematic. How can platforms reliably and proactively identify these “trust and safety” issues?
Machine learning has proved an attractive approach in the academic literature, leading to large bodies of scholarship on misinformation detection [154], toxic speech classification [357], and other core trust and safety challenges (e.g., [358]). The conceptual appeal of machine learning is that it could address the massive scale of user-generated content on large platforms and the capacity constraints of small platforms. Recent work claims impressive performance statistics: In the literature review that we conduct for this work, among publications that report performance metrics, about 70% of papers report over 80% accuracy on at least one detection task; some of these works report near-perfect performance [43, 44].
In the past year, news items from major tech companies have tempered these expectations. In June of 2023, OpenAI announced the deployment of a content moderation system based on GPT-4 to detect problematic content online; in the same press release, OpenAI’s head of safety systems admitted that human advisors would still need to “adjudicate borderline cases” [355]. In October of 2023, in a Bluesky post commenting on Twitter’s user-driven Community Notes program, Twitter’s former head of trust and safety stated that large-scale automated detection of misinformation remains a hard problem, and that no generalizable automated solutions are currently available [380]. These disclosures accord with our observation that, in practice, trust and safety functions at online services remain heavily manual: driven by user reports and carried out by human moderators.
In this work, we investigate the disconnect between scholarship and practice in applications of machine learning to trust and safety problems. Our project is inspired by recent research that has identified shortcomings in machine learning applications for many problem domains, including information security [331, 336]. We use misinformation detection as a case study for trust and safety problems because the topic has recently generated a rich literature with diverse methods and claimed successes. Misinformation detection has substantive complexities that are common for trust and safety problems: linguistic and cultural nuance, sensitivity to context, and rapidly evolving circumstances.
@@ 2 MOTIVATION
The detection of misinformation and “influence operations” (IOs) is a topic of convergent interest for security, social science, and AI/ML researchers. Methods developed in recent years for the detection of influence operations (e.g., astroturfing, coordinated misinformation campaigns) resemble techniques previously employed by security researchers for the detection of botnets, advanced persistent threats (APTs), and malware [378, 381, 359]. As such, we believe that the security community is uniquely well-positioned to shepherd the responsible development of misinformation detection methods. In addition, misinformation research stands to benefit a good deal from the formal structure of security research methodology: In a survey study from 2022 conducted by Mirza et al., human fact-checkers and journalists expressed interest in adopting rigorous threat modeling practices similar to those found in academic security literature [360].
@@ 3 PRELIMINARIES
Definitions. In the absence of agreed-upon definitions of misinformation, disinformation, malinformation, and influence operations, we refrain from advancing singular definitions of the same [^1] While certain taxonomies [154, 153, 155] provide relative definitions of these terms—distinguishing misinformation from disinformation, for instance, via the presence or absence of intent—we emphasize that, to the best of our knowledge, there is no robust invariant for identifying a statement as (mis)informative on the basis of semantics alone.. Instead, for each paper we review, we consider the paper authors’ working definitions of these terms (if such definitions are provided) and evaluate model performance with respect to the definitions set forth [258, 26, 25, 27, 23]. We limit our analysis to text-based English-language misinformation. [^2] We acknowledge that the nature of misinformation narratives and misinformation spread varies with language and geography. We note that disinformation is commonly used to refer to incorrect information written with the intent to deceive, while misinformation refers to incorrect information in general; for the sake of completeness and concision, we refer to intentionally and unintentionally false information as misinformation for this work unless otherwise qualified. We occasionally use false rumors, false news, and false information interchangeably, but avoid fake news, a politically charged term [424]. We use “influence operations” (IOs) or “coordinated campaigns” in our discussion of collaborative attempts to disseminate misinformation across networks [425].
@@
Feasibility of detection. Most of the work we review presupposes that automated detection of misinformation is possible at all. We note that this is, in itself, a strong assumption to make: Published work in psychology, linguistics, and philosophy of language has previously called into question the feasibility of using semantic and syntactic features to determine the veracity of text statements and the usefulness of binary true/false classifications of information, particularly in the absence of unified definitions of misinformation, disinformation, and information [354, 201]. On the other hand, some researchers maintain that misinformative texts are characterized by indelible “fingerprints” that distinguish them from non-misinformative texts: for instance, emotional language and reduced lexical diversity [177, 327]. We use this debate to motivate our discussion of non-textual feature sets for misinformation detection.
@@
Taxonomies. Through an iterative process of reading and inductive coding of our papers, we develop a taxonomy of five “information scopes.” As we read each paper in our corpus, we noted the operative unit of misinformation detection for the classifier or method described—the smallest semantic or organizational unit of information that the method attempted to classify as true or false—and sorted each paper into emergent categories. Our final taxonomy comprises five information scopes: claims, news articles, social media accounts, networks, and websites. We consider these scopes in the context of online services—for instance, news articles and websites whose links might be posted to a social media platform. [^3] We label certain works as members of multiple scopes when they make classification decisions about more than one type of detection unit: works targeting social media posts, for instance, are generally categorized as falling within ( C and U or ( C and N) in Table I.
@@
    Networks. The set of users and interactions between these users as represented by a networked graph of behaviors.
@@
Paper organization. In Section 2, we situate this project within existing security literature and academic literature that critiques machine learning-driven methods. We motivate our choice of automated misinformation detection as a representative case study and highlight the particular relevance of this problem to the security community. In Section 3, we present the information taxonomies developed from our inductive coding of papers. We discuss findings from our reading and coding of papers in Section 4. Our systematization of literature progresses along two axes: 1) from unimodal to increasingly multi-modal methods, and 2) in order of operations of method development. Specifically, we discuss issues pertaining to method-target fit, dataset curation and feature selection, model selection, and method evaluation (RQ1, RQ2). In Section 5, we illustrate the issues identified in our literature review with a series of replication studies, with a focus on reproducibility and generalizability of results (RQ3, RQ4). In Section 6, we conclude with 1) a discussion of findings from our literature review and replication studies and 2) generalizable recommendations for evaluating ML-driven interventions for trust and safety.
@@ 4 Systematization of Literature
To ensure that security-oriented approaches to detection were represented in our corpus, we conducted a separate snowball sampling search for work published in security venues: Using the same keywords listed in the previous subsection, we oversampled publications from four prominent [^6] A*-rated venues, as determined by CORE [255] security research venues (USENIX Security, IEEE S&P, NDSS, and ACM CCS). This sampling process resulted in the addition of 29 additional works, most of which address the detection of accounts and networks that spread misinformative content (e.g., botnets and trolls); as such, most of these works are categorized as network- or account-scoped detection methods. Our final corpus comprises 248 papers published between 2009 and 2023, inclusive.
@@
“Full” and “focus” paper corpora. We conduct our literature survey at two different levels of granularity. For all papers in the full corpus (248 works), we note detection targets and model evaluation (steps 1 and 5 in Section 3). For a subset of these (87), we perform deep-coding of methods: we note dataset design choices, model selection, and feature sets (resp. steps 2, 3, and 4 in Section 3). [^7] We take inspiration from [2, 4, 3], who perform a similar deep-coding of a subset of their whole-paper corpus. A summary of this deep-coding is available in Table I. Our motivation for developing this focus set is practical: A good many of the works we review do not include in-depth discussions of model choice, weights, and datasets. As such, each per-scope discussion is book-ended by comprehensive overviews of the full paper corpus, with more focused analyses of method design details in between.
@@
Corpus curation. As a first-pass relevancy check, we used automated keyword matching to confirm that all collected papers did, in fact, address misinformation and automated fact-checking methods; we read the abstracts of papers surfaced by this check to confirm relevance and fit. We then annotated research papers that passed this check in accordance with our codebook (see SI). One reader made three separate coding passes over the corpus, varying the objectives of her annotation on each pass: In the first pass, she specifically sought to identify taxonomies for classification and detection; on the second, she noted actual versus stated detection targets and approaches to evaluation; on the final pass, she noted method design approaches and errors.
@@
In- and out-of-scope work. We consider a number of security-oriented approaches to misinformation detection in this work. Sybil and botnet detection methods that specifically target influence operations online are categorized as account- and network-scoped detection methods in our corpus. Commercial approaches to misinformation detection are in-scope for this project. In view of data accessibility issues, however, we are unable to provide an in-depth analysis of commercial methodologies in our main literature review, and instead include a market survey of commercial fact-checking providers in our online supplement. [^8] https://anonymous.4open.science/r/sok_misinformation-B297/ (In general, commercial vendors do not make code and training data publicly available.) LLM-powered detection is in-scope for this work. Accessibility to code bases for LLMs is similarly limited (this has been discussed at length in the popular press [393, 394]), and disallows extensive testing and evaluation by researchers. As such, while we do briefly discuss LLM-powered detection approaches in our supplement, we defer a more extensive discussion to future work.
@@
Notation. In these sections, we denote information scope and pipeline steps with circled icons ( A, 1) and subtaxa within those categories with lowercase bolded Roman numerals (iv). We note that, while targets are generally unique to each information scope (and are referenced by an alphanumeric label (e.g., (A.i)), critiques of steps 2- 5 are not scope-specific, and are always denoted by Arabic-Roman numeral pairs (e.g., (2.i)). We highlight each subtaxon in-text, with a different color corresponding to each development step. (The full taxonomy is in Section 7.) A summary of findings for each scope is available in the Takeaways box at the end of each subsection.
@@
Network-based
Author-based
Infrastructural
@@
1. Ajao et al. [278] C A Sentiment (A.i) PHEME [146] LSTM, DT, RF, SVM 0.86 (Acc.)
@@
2. Abulldah-All-Tanvir et al. [210] C A N Content (C.i) Twitter (API) NB, RNN, LSTM, SVM, Logit 0.89 (Acc.)
@@
3. Bhutani et al. [236] C A Content (C.i); sentiment (A.i) Twitter (API), PolitiFact [299] Naive Bayes, RF 0.60 (AUC)
4. Bozarth et al. [20] C Contents (C.i) PolitiFact [299], Daily Dot, Zimdars, MBFC LDA n/a
5. Ciampaglia et al. [22] C N Shortest path search (C.ii) DBpedia kNN, RF 0.97 (AUC)
6. Cui et al. [277] C A U Content (C.i); sentiment (A.i); user response (U.ii) PolitiFact[299], GossipCop[343] KNN, SVM, CSI [37], RMSprop 0.82 (F1)
7. Debnath et al. [84] C Content (C.i) LIAR [35] CNN 0.27 (Acc.)
8. Dey et al. [241] C A Content (C.i); sentiment (A.i) Twitter (API) Clustering (kNN) 0.67 (Acc.)
9. Galitsky et al. [215] C Content (C.i) Amazon reviews Parse thicket 0.81 (Prec.)
10. Glockner et al. [142] C Content (C.i) PolitiFact[299], Snopes[298], MultiFC CNN, DNN 0.58 (Acc.)
11. Gordon et al. [110] C A Content (C.i); source rep (U.i) Credibility-Factors2020 SVD 0.63 (Acc.)
12. Gupta et al. [294] C A N Content (C.i); stance (A.iii) Twitter (API) SVM 0.60 (Agreement)
13. Hassan et al. [130] C Content (C.i); checkability (C.iii) NBA, weather datasets Frequency n/a
14. Jain et al. [291] C Content (C.i); sentiment (A.i) Twitter (API) Gensim/TextBlob 0.77 (Acc.)
15. Jiang et al. [32] C Content (C.i); ling./syntax (C.ii) PolitiFact[299], Snopes[298] SVM 0.81 (Acc.)
16. Karimi et al. [260] C Content (C.i) LIAR[35] LSTM, CNN 0.39 (Acc.)
17. Kartal et al. [108] C Content (C.i); checkability (C.iii) Check That! dataset Logit, SVM, RF 0.26 (MAP)
18. Kou et al. [128] C Content (C.i) CoAID, CONSTRAINT Knowledge graph 0.90 (Acc.)
19. Paudel et al. [185] N A Keyword detection (A.ii) Abilov et al. dataset, Twitter (API) AdaRank, ListNet, RF 0.79 (MAP)
20. Popat et al. [90] C Content (C.i) PolitiFact[299], Snopes[298], NewsTrust biLSTM, CNN 0.88 (AUC)
21. Shiralkar et al. [56] C N KG search (C.ii) DBpedia Knowledge graph 1.00 (AUC)
22. Shu et al. [207] C U Word encoding (C.i); user response (U.ii) GossipCop [343], PolitiFact [299] RNN/RMSprop, CSI [37], LSTM, CNN 0.93 (F1)
23. Tian et al. [106] C U Content (C.i); user response (U.ii) Twitter15, Twitter16 CNN-biLSTM 0.82 (F1)
24. Zhang et al. [382] C U Content (C.i); user response (U.ii) RumourEval, PHEME[146] biLSTM, Multitask, SVM, CNN, 0.89 (Acc.)
@@
1. Afroz et al. [212] A Content (C.i); syntax (A.i) Brennan-Greenstadt SVM, J48 Decision Trees 0.97 (F1)
@@
2. Ahmed et al. [214] A Syntax (A.i) Twitter, Kaggle, Horne and Adali [24] SVM 0.92 (Acc.)
@@
3. Bourgonje et al. [55] A Stance (A.ii) Fake News Challenge Data Logit 0.90 (Acc.)
4. Brasoveanu et al. [68] A C Sentiment (A.i); keywords (A.ii) LIAR[35] CNN, LSTM, CN 0.64 (Acc.)
@@
5. Della Vedova et al. [58] A N Content (C.i); virality (N.iv) FakeNewsNet, Buzzfeed Logit 0.82 (Acc.)
6. Horne et al. [24] A Syntax (A.i), headline (A.ii) Buzzfeed, Burfoot & Baldwin SVM 0.78 (Acc.)
7. Jabiyev et al. [118] A W Topic detection (A.ii); site cred. (W.iv) Snopes[298], FactCheck, PolitiFact[299] SVM, DT, RF 0.87 (Acc.)
8. Jadhav et al. [238] A Content (C.i); syntax (A.i) LIAR [35] DSSM/RNN 0.99 (Acc.)
9. Jin et al. [216] A N U (A.ii); (N.i); suspicious accounts (U.i) Tweets; articles n/a 0.87 (Prec.)
10. Kapusta et al. [88] A Sentiment & word freq. (A.i) MBFC and custom n/a n/a
11. Kumar et al. [29] A W U (A.i); UI (W.iii); Author cred. (U.i) 20K Wiki Hoaxes Random forest 0.87 (AUC)
12. Magdy et al. [18] A Content (C.i) NYT Corpus [224], 100 Wikis Pattern recog. 0.99 (Recall)
13. Monti et al. [315] A N U (A.ii); (N.i); suspicious accounts (U.i) Tweets; articles RNN/CNN 0.927 (AUC)
14. Nasir et al. [141] A Syntax (A.i) ISOT [312]; FAKES [313] RNN/CNN 0.99 (Acc.)
15. Perez-Rosas et al. [63] A Syntax (A.i) FakeNewsAMT; Celebrity SVM 0.74 (Acc.)
16. Potthast et al. [17] A Syntax, sentiment, readability (A.i) Buzzfeed-Webis Bag-of-words 0.46 (F1)
17. Reis et al. [286] A Syntax (A.i); contents (C.i); source rep. (U.i); timing (N.ii); URL (W.i) BuzzFeed GBM 0.85 (AUC)
18. Rubin et al. [292] A Syntax (A.i) AMT Clustering 0.67 (Agreement)
19. Ruchansky et al. [37] A U (A.i); Responses (A.ii); acct metadata (U.i) Twitter/Weibo posts RNN/LSTM 0.95 (Acc.)
20. Santos et al. [19] A Readability (A.i) Fake.Br corpus SVM 0.92 (Acc.)
21. Silva et al. [111] A N Topic detection (A.ii); propagation (N.i) PolitiFact[299], GossipCop[343], CoAID Clustering 0.88 (Acc.)
21. Singh et al. [203] A Syntax (A.i) Kaggle Fake News SVM 0.87 (Acc.)
22. Uppal et al. [102] A Discourse structure (A.i) Buzzfeed, PolitiFact[299] GRU, Dependency tree 0.74 (Acc.)
@@
1. Cao et al. [101] U N Acct. cred. (U.i); prop. (N.i) Tuenti social network Louvain clustering 0.90+ (TP)
@@
2. Danezis et al. [373] U N Acct. cred. (U.i); prop. (N.i) LiveJournal data Bayesian inf. n/a*
3. Ezzeddine et al. [366] U Acct. behaviors (U.ii) DATA LSTM 0.91 (AUC)
@@
4. Hamdi et al. [91] U N Account metadata (U.i); prop. (N.i) CREDBANK LDA, Bayes, Logit, SVM 0.99 (AUC)
@@
5. Helmstetter et al. [5] U N A Acct metadata (U.i); post sharing data (U.ii) Public site cred. lists SVM, NB, DT, RF 0.936 (F1)
6. Jain et al. [291] U A Acct metadata (U.i); topic det. (A.ii) Twitter (API) Gensim, TextBlob 0.77 (Acc.)
7. Leonardi et al. [38] U N A Acct metadata (U.i); prop. (N.i) CoAID RF 0.81 (F1)
8. Saeed et al. [361] U User behavior (U.ii) Reddit Pushshift; Reddit IRA trolls list RF 0.98 (Acc.)
9. Sansonetti et al. [384] U C Acct metadata (U.i); acct activity (U.ii) PolitiFact [299], Twitter (API) LSTM-CNN, SVM, KNN 0.92 (Acc.)
10. Santia et al. [21] U A Source rep. (U.i); user response (U.ii); syntax (A.i) BuzzFeed SVM, RF, DT, NB 0.77 (Prec.)
11. Shu et al. [226] U N A User behavior (U.ii); prop. (N.i) syntax (A.i) Buzzfeed, PolitiFact Gibbs sampling 0.85+ (Acc.)
12. Vargas et al. [376] N A Prop. (N.i); topic det. (A.ii) Twitter (API) RF 0.98 (F1)
13. Wang et al. [191] U N User behavior (U.ii) Renren data SVM 0.99 (Acc.)
14. Yu et al. [200] U N User behavior (U.i); prop. (N.i) LiveJournal, Friendster, DBLP accounts Random route n/a*
15. Yuan et al. [375] U N Acct metadata (U.i); timing (N.ii) WeChat data Clustering 0.90+ (Prec.)
16. Zhang et al. [257] U N User behavior (U.ii); prop. (N.i) Twitter, Slashdot, Epinion Graph cut n/a
17. Zhou et al. [72] U N User suscept. (U.i); prop (N.i) PolitiFact [299], BuzzFeed SVM, KNN, NB, DT, RF 0.93 (Acc.)
@@
1. Alizadeh et al. [370] N A U Propagation (N.i); syntax (A.i); acct metadata (U.i) Twitter (API), Reddit IRA troll list RF 0.70+ (F1)

### v4

# SoK: Machine Learning for Misinformation Detection

@@ 1 INTRODUCTION
Online services face a daunting task: There is an unceasing deluge of user-generated content, on the order of hundreds of thousands of posts per minute on popular social media platforms [dewan_2015]. Some of that content is false, hateful, harassing, extremist, or otherwise problematic. How can platforms reliably and proactively identify these “trust and safety” issues?
Machine learning has proved an attractive approach in the academic literature, leading to large bodies of scholarship on misinformation detection [alam_survey_2021], toxic speech classification [andročec_2021], and other core trust and safety challenges (e.g., [lee_2020]). The conceptual appeal of machine learning is that it could address the massive scale of user-generated content on large platforms and the capacity constraints of small platforms. Recent work claims impressive performance statistics: In the literature review that we conduct for this work, among publications that report performance metrics, about 70% of papers report over 80% accuracy on at least one detection task; some of these works report near-perfect performance [baly_predicting_nodate, dhamani_using_2019].
In recent years, news items from major tech companies have tempered these expectations. In October of 2023, in a Bluesky post commenting on Twitter’s user-driven Community Notes program, Twitter’s former head of trust and safety stated that large-scale automated detection of misinformation remains a hard problem, and that no generalizable automated solutions are currently available [roth_2023]. In January of 2025, Meta announced that it would terminate its third-party fact-checking program in favor of a Community Notes-like system of user-driven content moderation [chan_meta_2025]. These disclosures accord with our observation that, in practice, trust and safety functions at online services remain heavily manual: driven by user reports and carried out by human moderators.
In this work, we investigate the disconnect between scholarship and practice in applications of machine learning to trust and safety problems. Our project is inspired by recent research that has identified shortcomings in machine learning applications for many problem domains, including information security [arp_ml_2022, Jacobs2022]. We use misinformation detection as a case study for trust and safety problems because the topic has recently generated a rich literature with diverse methods and claimed successes. Misinformation detection has substantive complexities that are common for trust and safety problems: linguistic and cultural nuance, sensitivity to context, and rapidly evolving circumstances.
@@ 2 MOTIVATION
The detection of misinformation and “influence operations” (IOs) is a topic of convergent interest for security, social science, and AI/ML researchers. Methods developed in recent years for the detection of influence operations (e.g., astroturfing, coordinated misinformation campaigns) resemble techniques previously employed by security researchers for the detection of botnets, advanced persistent threats (APTs), and malware [milajerdi2019holmes, king2006subvirt, stone2009botnet]. As such, we believe that the security community is uniquely well-positioned to shepherd the responsible development of misinformation detection methods. In addition, misinformation research stands to benefit a good deal from the formal structure of security research methodology: In a survey study from 2022 conducted by Mirza et al., human fact-checkers and journalists expressed interest in adopting rigorous threat modeling practices similar to those found in academic security literature [mirza2023tactics].
@@ 3 PRELIMINARIES
Definitions. In the absence of agreed-upon definitions of misinformation, disinformation, malinformation, and influence operations, we refrain from advancing singular definitions of the same [^1] While certain taxonomies [alam_survey_2021, zhou_survey_2020, oshikawa_survey_2018] provide relative definitions of these terms—distinguishing misinformation from disinformation, for instance, via the presence or absence of intent—we emphasize that, to the best of our knowledge, there is no robust invariant for identifying a statement as (mis)informative on the basis of semantics alone.. Instead, for each paper we review, we consider the paper authors’ working definitions of these terms (if such definitions are provided) and evaluate model performance with respect to the definitions set forth [wu_misinformation_2019, gelfert_fake_2018, haiden_definitional_2018, klein_fake_2017, tandoc_defining_2018]. We limit our analysis to text-based English-language misinformation. [^2] We acknowledge that the nature of misinformation narratives and misinformation spread varies with language and geography. We note that disinformation is commonly used to refer to incorrect information written with the intent to deceive, while misinformation refers to incorrect information in general; for the sake of completeness and concision, we refer to intentionally and unintentionally false information as misinformation for this work unless otherwise qualified. We occasionally use false rumors, false news, and false information interchangeably, but avoid fake news, a politically charged term [habgood_coote_2018]. We use “influence operations” (IOs) or “coordinated campaigns” in our discussion of collaborative attempts to disseminate misinformation across networks [jackson_2023].
@@
Feasibility of detection. Most of the work we review presupposes that automated detection of misinformation is possible at all. We note that this is, in itself, a strong assumption to make: Published work in psychology, linguistics, and philosophy of language has previously called into question the feasibility of using semantic and syntactic features to determine the veracity of text statements and the usefulness of binary true/false classifications of information, particularly in the absence of unified definitions of misinformation, disinformation, and information [soe_2017, vrij_pitfalls_2008]. On the other hand, some researchers maintain that misinformative texts are characterized by indelible “fingerprints” that distinguish them from non-misinformative texts: for instance, emotional language and reduced lexical diversity [amado_undeutsch_2015, carrasco_2022]. This debate motivated our discussion of non-textual feature sets for detection.
@@
Taxonomies. Through an iterative process of reading and inductive coding of our papers, we develop a taxonomy of five “information scopes.” As we read each paper in our corpus, we noted the operative unit of misinformation detection for the classifier or method described—the smallest semantic or organizational unit of information that the method attempted to classify as true or false—and sorted each paper into emergent categories. Our final taxonomy comprises five information scopes: claims, news articles, social media accounts, networks, and websites. We consider these scopes in the context of online services—for instance, news articles and websites whose links might be posted to a social media platform. [^3] We label certain works as members of multiple scopes when they make classification decisions about more than one type of detection unit: works targeting social media posts, for instance, are generally categorized as falling within ( C and U) or ( C and N) in Table 4.
@@
    Networks. A set of users and interactions between these users as represented by a social graph.
@@
Paper organization. In Section 2, we situate this project within existing security literature and academic literature that critiques machine learning-driven methods. We motivate our choice of automated misinformation detection as a representative case study and highlight the particular relevance of this problem to the security community. In Section 3, we present the information taxonomies developed from our inductive coding of papers. We discuss findings from our reading and coding of papers in Section 4. Our systematization of literature progresses along two axes: 1) from unimodal to increasingly multi-modal methods, and 2) in order of operations of method development. Specifically, we discuss issues pertaining to method-target fit, dataset curation and feature selection, model selection, and method evaluation (RQ1, RQ2). In Section 5, we illustrate the issues identified in our literature review with a series of replication studies, with a focus on reproducibility and generalizability of results (RQ3, RQ4). In Section 6, we conclude with 1) a discussion of findings from our literature review and replication studies and 2) recommendations for evaluating ML-driven interventions for trust and safety.
@@ 4 Systematization of Literature
“Full” and “focus” paper corpora. We conduct our literature survey at two different levels of granularity. For all papers in the full corpus (248 works), we note detection targets and model evaluation (steps 1 and 5 in Section 3). For a subset of these (87), we perform deep-coding of methods: we note dataset design choices, model selection, and feature sets (resp. steps 2, 3, and 4 in Section 3). [^6] We take inspiration from [wei_sociodem_2024, scheffler_e2ee_2023, warford_ux_2021], who perform a similar deep-coding of a subset of their whole-paper corpus. A partial summary of this deep-coding is available in Table 2. Our motivation for developing this focus set is practical: Many of the works we review do not include in-depth discussions of model choice, weights, and datasets. As such, each per-scope discussion is book-ended by overviews of the full paper corpus, with focused analyses of method design details in between.
Targets C Claims: i. Content-based detection via distance calculations on semantic embeddings; ii. Content-based detection via search on knowledge graph topology; iii. “Checkability” or “checkworthiness.” A Articles: i. Syntactic and stylometric signals, including genre and sentiment; ii. Topic-aware detection of stance and relevance to known rumoring topics. U Users: i. Account metadata (bios, images, account age); ii. Single account behaviors (comments on posts, published posts). N Networks: i. Propagation patterns across social graphs; ii. Timestamped records of user-user interactions. W Websites: i. Text and URL/domain semantics; ii. Site visitor demographics; iii. Suspicious UI elements; iv. Hosting infrastructure (DNS certificate, site age). Datasets i. Dataset age; ii. Evidence of leakage (temporal leakage, feature leakage?); iii. Data dependencies (author, source, style); iv. Availability of information required to reproduce or reconstruct similar datasets (was non-public information required to produce ground-truth training sets?); v. Availability of original data. Models i. Distance calculations on semantic embeddings; ii. “Traditional” ML (SVM, RF, DT); iii. Deep learning (CNN, LSTM, GRU); iv. Graph cut algorithms; v. Stacked ensemble classifiers; vi. Graph clustering algorithms. Features i. Textual; ii. Network-based; iii. Author-, user- or source-based; iv. Infrastructural. Evaluation i. Testing in real time; ii. Generalizability of approach; iii. Evasion-resistance; iv. Robustness to distributional shifts in training or test data; v. False-positive/false-negative rates.
Table 1: Taxonomy of targets, models, features, and evaluation codes. Alphanumeric abbreviations (e.g., “C.i”) are referenced throughout this work.
@@
Corpus curation and coding. As a first-pass relevancy check, we used automated keyword matching to confirm that all collected papers did, in fact, address misinformation and automated fact-checking methods; we read the abstracts of papers surfaced by this check to confirm relevance and fit. We then annotated research papers that passed this check in accordance with our codebook (see SI). One reader made three separate coding passes over the corpus, varying the objectives of her annotation on each pass: In the first pass, she specifically sought to identify taxonomies for classification and detection; on the second, she noted actual versus stated detection targets and approaches to evaluation; on the final pass, she noted method design approaches and errors. Two coders read and independently coded a random subset (30 papers) of the full corpus. Fleiss’s kappa for this subset was > 0.80, indicating strong agreement.
@@
In- and out-of-scope work. We consider a number of security-oriented approaches to misinformation detection in this work. Sybil and botnet detection methods that specifically target influence operations online are categorized as account- and network-scoped detection methods in our corpus. Commercial approaches to misinformation detection are in-scope for this project. In view of data accessibility issues, however, we are unable to provide an in-depth analysis of commercial methodologies in our main literature review, and instead include a market survey of commercial fact-checking providers in Appendix A. (In general, commercial vendors do not make code and training data publicly available.) LLM-powered detection is in-scope for this work. Accessibility to code bases for LLMs is similarly limited (this has been discussed at length in the popular press [mollman_2023, wong_2024]), and disallows testing and evaluation by researchers. As such, while we do briefly discuss LLM-powered detection approaches in our supplement, we defer a more extensive discussion to future work.
@@
Notation. In these sections, we denote information scope and pipeline steps with circled icons ( A, 1) and subtaxa within those categories with lowercase bolded Roman numerals (iv). We note that, while targets are generally unique to each information scope (and are referenced by an alphanumeric label (e.g., (A.i)), critiques of steps 2- 5 are not scope-specific, and are always denoted by Arabic-Roman numeral pairs (e.g., (2.i)). We highlight each subtaxon in-text, with a different color corresponding to each development step. (The full taxonomy is in Table 1.) A summary of findings for each scope is available in the starred **Takeaways summary at the end of each subsection.
2
Datasets. Claim-scoped papers in our corpus that propose to perform fact verification (2.i) rely on outdated existing datasets of labeled statements in order to establish ground truth. LIAR [wang_liar_2017] [^7] We note that the LIAR dataset is actually a collection of 27.8K labeled PolitiFact statements—so, in a sense, PolitiFact is the dominant data source. and PolitiFact [politifact] were the most popular datasets among claim-scoped works, and, taken together, were used by approximately half of all papers within this scope [wang_learning_2019, soprano_many_2021]. LIAR is a static political news dataset that was published in 2017; on average, papers that cite LIAR were published two years after LIAR’s release. Topic detection and word frequency models trained on LIAR are likely ineffective in contemporary fact-checking contexts, and for non-political subject matter [debnath_hierarchical_2020, rasool_multi-label_2019]. [^8] Choice of ground truth site or labeled dataset can significantly influence the outcome of analysis: Bozarth et al. found that perceived prevalence of misinformation in a corpus of 2016 election news varied from 2% to 40%, depending on choice of ground-truth reference website [bozarth-truth-2020]. Additionally, misinformation taxonomies across reference sites are inconsistent: PolitiFact employs a six-point labeling scale (pants-on-fire; false; barely-true; half-true; mostly-true; true), FEVER employs a three-point scale (supported; refuted; notenoughinfo), and GossipCop employs an eleven-point scale (ratings from 0 to 10). This divergence is likely a symptom of definitional issues (Section 3).
3
Model selection. Model choice follows target choice for claim-scoped methods: for methods that pre-construct knowledge graphs or other reference databases, models perform some form of (3.i) shortest-path search on the KG topology [shiralkar_finding_2017, ciampaglia-wiki-2015], and approximate logical inference via transitive closure on graph edges. For methods that perform information retrieval at query time—e.g., to match corroborating sources to a claim to be checked—model training generally follows conversion of the text statement to a bag-of-words or TF-IDF embedding; choice of model is highly variable [bhutani_fake_2019, shu_defend_2019], and does not appear to be predictive of performance. For methods that perform detection of misinformative posts on social media, (3.v) stacked ensemble classifiers are a common approach to incorporating multiple feature modalities.
@@
Network
Author
Infra.
@@
1. Ajao et al. [ajao_sentiment_2019] C A Sentiment (A.i) PHEME [pheme_2018] LSTM, DT, RF, SVM 0.86 (Acc.)
@@
2. Abulldah-All-Tanvir et al. [abdullah-all-tanvir_detecting_2019] C A N Content (C.i) Twitter (API) NB, RNN, LSTM, SVM, Logit 0.89 (Acc.)
@@
3. Bhutani et al. [bhutani_fake_2019] C A Content (C.i); sentiment (A.i) Twitter (API), PolitiFact [politifact] Naive Bayes, RF 0.60 (AUC)
4. Bozarth et al. [bozarth-truth-2020] C Contents (C.i) PolitiFact [politifact], Daily Dot, Zimdars, MBFC LDA n/a
5. Ciampaglia et al. [ciampaglia-wiki-2015] C N Shortest path search (C.ii) DBpedia kNN, RF 0.97 (AUC)
@@
1. Afroz et al. [afroz_detecting_2012] A Content (C.i); syntax (A.i) Brennan-Greenstadt SVM, J48 Decision Trees 0.97 (F1)
@@
2. Ahmed et al. [ahmed_detecting_2018] A Syntax (A.i) Twitter, Kaggle, Horne and Adali [horne_this_2017] SVM 0.92 (Acc.)
@@
3. Bourgonje et al. [bourgonje_clickbait_2017] A Stance (A.ii) Fake News Challenge Data Logit 0.90 (Acc.)
4. Brasoveanu et al. [brasoveanu_semantic_2019] A C Sentiment (A.i); keywords (A.ii) LIAR[wang_liar_2017] CNN, LSTM, CN 0.64 (Acc.)
@@
5. Della Vedova et al. [della_vedova_automatic_2018] A N Content (C.i); virality (N.iv) FakeNewsNet, Buzzfeed Logit 0.82 (Acc.)
@@
1. Cao et al. [cao_aiding_2012] U N Acct. cred. (U.i); prop. (N.i) Tuenti social network Louvain clustering 0.90+ (TP)
@@
2. Danezis et al. [danezis2009sybilinfer] U N Acct. cred. (U.i); prop. (N.i) LiveJournal data Bayesian inf. n/a*
3. Ezzeddine et al. [ezzeddine2023exposing] U Acct. behaviors (U.ii) DATA LSTM 0.91 (AUC)
@@
4. Hamdi et al. [hamdi_hybrid_2020] U N Account metadata (U.i); prop. (N.i) CREDBANK LDA, Bayes, Logit, SVM 0.99 (AUC)
@@
5. Helmstetter et al. [helmstetter-supervised-2018] U N A Acct metadata (U.i); post sharing data (U.ii) Public site cred. lists SVM, NB, DT, RF 0.936 (F1)
@@
1. Alizadeh et al. [alizadeh2020content] N A U Propagation (N.i); syntax (A.i); acct metadata (U.i) Twitter (API), Reddit IRA troll list RF 0.70+ (F1)